How many electrons are there in an atom if. a) n=2 , s=1/2 & b) n=4, l=2

Grade:11

2 Answers

Ashish
6 years ago
a)for n=2 , l=1,0 m= 3and m=1 for l=1 and l=0 respectively so for m=3 total electrons are 6 as each orbital has 2 electrons out of which 3 has s= 1/2 and for m=1 total electrons is 2 and 1 has s = 1/2 so ....Total electrons for given values is 4.b) for n=4 and l=2 m=5 and each orbital has 2 electrons so 5 orbitals has 10 electrons.
Deepak Kumar
6 years ago
(a). For n=2 and s = 1/2 ,Actually for n=2 , l = 0 , 1 i.e. 2s and 2p orbitalAnd in each orbital there will be one electron for s =1/2So, taking this into account, answer will be 4 electrons.(b). For n= 4 and l = 2 ,l =2 corresponds to d orbital .So we have 4d from above given condition.And for l=2 , m has 5 values as -2,-1,0,1,2.So there are 5 orbitals and each orbital has 2 electrons .So, the answer will be 5*2 = 10 electrons